OpenStudy (anonymous):

(3x^-5)(-6x^7y^8)(1/3y^-4) Simplify please

OpenStudy (anonymous):

\[-6x^2y^4\]

OpenStudy (anonymous):

well, when I did it, I got -8x^2y^4/3 How did you get your answer?

OpenStudy (anonymous):

oh never mind, I put -8 instead of -18.. now I feel stupid

OpenStudy (anonymous):

multiply all the constants in front, add the respective exponents for x and y 3*-6*(1/3)=-6 x^-5*x^7=x^2 y^8*y^-4=y^4 put it all together 6x^2y^4

OpenStudy (anonymous):

-*

OpenStudy (anonymous):

Gracias

OpenStudy (anonymous):

(3 x^-5) (-6 x^7 y^8) (1/3 y^-4)\[\left(\frac{3}{x^5}\right)\left(-6 x^7 y^8\right)\left(\frac{1}{3 y^4}\right) \]\[-6 x^2 y^4 \]